# TW-CRL: Time-Weighted Contrastive Reward Learning for Efficient Inverse Reinforcement Learning

### v1

@@ 4.2.3 TW-CRL: Complete Algorithm Pipeline
(d) PandaReach
(e) PandaPush
Figure 2: Training and success rate curves on benchmarks. The solid curves represent the mean, and the shaded regions indicate the standard deviation over five runs. The first row shows the average return curves for each environment, while the second row shows the success rate curves for each environment.
Table 1: Average final return and success rate across benchmarks. The first sub-rows show the average return, and the second sub-rows show the success rate. The best performance is marked in bold, and ± represents the standard deviation over five runs.
Environment TW-CRL SASR GAIL AIRL MaxEnt TD3
U-Maze 241.57 ± 43.22 192.72 ± 78.15 138.64 ± 20.25 208.21 ± 17.83 161.28 ± 41.38 226.67 ± 45.53
0.91 ± 0.18 0.76 ± 0.30 0.64 ± 0.19 0.80 ± 0.15 0.68 ± 0.31 0.86 ± 0.18
TraptMaze-v1 241.85 ± 8.80 196.28 ± 32.89 159.06 ± 25.53 160.22 ± 11.25 129.70 ± 18.50 127.83 ± 26.67
0.94 ± 0.07 0.82 ± 0.18 0.78 ± 0.09 0.72 ± 0.08 0.46 ± 0.19 0.46 ± 0.20
TrapMaze-v2 207.64 ± 11.67 63.68 ± 26.71 113.94 ± 50.26 48.95 ± 19.92 77.89 ± 30.10 97.00 ± 35.04
0.89 ± 0.08 0.23 ± 0.14 0.52 ± 0.18 0.21 ± 0.15 0.31 ± 0.15 0.33 ± 0.09
PandaReach-v3 -5.98 ± 0.73 -47.41 ± 0.51 -46.53 ± 2.28 -10.76 ± 5.93 -6.01 ± 2.30 -37.58 ± 7.49
1.00 ± 0.00 0.03 ± 0.51 0.10 ± 0.14 0.86 ± 0.19 1.00 ± 0.00 0.33 ± 0.17
PandaPush-v3 -34.89 ± 7.59 -96.77 ± 2.34 -99.00 ± 1.40 -97.39 ± 3.00 -96.45.00 ± 2.54 -73.96 ± 18.42
0.73 ± 0.05 0.10 ± 0.08 0.00 ± 0.00 0.07 ± 0.05 0.03 ± 0.05 0.37 ± 0.19
5 Experiments
In this section, we conduct experiments to evaluate the performance of TW-CRL, visualize and analyze the reward function in TrapMaze-v1, examine TW-CRL’s ability to generalize to various goal settings, and perform ablation studies.
5.1 Experimental Setup
We evaluate TW-CRL on five continuous control tasks, which include three navigation tasks and two robotic manipulation task. All tasks are episodic, with each episode having a fixed timestep limit with sparse reward. We repeat each experimental setting five times and report both average episodic return and the average success rate. Details regarding hyperparameters, network architectures, and additional training setups are provided in Appendix A and Appendix C.
Benchmarks We use a total of five environments. This includes four Point Maze environments (Fu et al., 2020; de Lazcano et al., 2024), including a classic U-Maze as well as two custom variants, TrapMaze-v1 and TrapMaze-v2, both of which contain unknown trap states and have randomly initialized goal states. Besides, we also use two robotic manipulation environment – PandaReach and PandaPush from panda-gym (Gallouédec et al., 2021; Plappert et al., 2018). The environments are illustrated in Figure 6.
Baselines To demonstrate the effectiveness of our approach, we compare TW-CRL with five baselines: SASR(Ma et al., 2024), GAIL(Ho & Ermon, 2016), AIRL(Fu et al., 2017), MaxEnt (Ziebart et al., 2008), and TD3(Fujimoto et al., 2018).
5.2 Comparison Evaulation
Figure 2 and Table 1 report the average episodic returns, their variances, and the average success rates across all tasks and baselines. We provide five successful demonstrations in the basic U-Maze environment, while TrapMaze-v1 and TrapMaze-v2 each receive 35 successful demonstrations, all ending in a randomly initialized goal located within a single designated grid area of the maze. During training, the same grid area is used for goal initialization, though the specific location within that grid varies each episode. In PandaReach and PandaPush, we offer 10 successful demonstrations.
(a) TW-CRL
(b) GAIL
(c) AIRL
(d) MaxEnt
(e) SASR
The results indicate that the performance of TW-CRL is better than or comparable to all baselines in terms of average episode return and success rate, with faster convergence and reduced variance. For example, in the TrapMaze-v1 environment, the average final return in our algorithm demonstrates enhancements of $23.2\%$, $52.0\%$, $50.9\%$, $86.5\%$, and $89.2\%$ over SASR, GAIL, AIRL, MaxEnt, and TD3, respectively.
5.3 Analysis of Learned Rewards
In this section, we analyze the learned reward functions across different methods using the TrapMaze environment to illustrate why TW-CRL outperforms other approaches. The simplified illustration of the TrapMaze-v1 environment map is shown in Figure 4(a), where the red area represents the goal states, the blue area represents the trap states, and the gray blocks indicate walls that the agent cannot pass through. We provide a total of 35 successful demonstrations in which agents follow the main path to the goal, as depicted in Figure 4(b). Figure 3 visualizes the reward functions learned by different methods throughout the training process.
The reward function visualizations reveal that GAIL, AIRL, and MaxEnt fail to recognize unseen trap states, and give high rewards only to states seen in successful demonstrations. As a result, these methods limit the agent to only imitating expert behavior without discovering better ways to reach the goal. On the other hand, SASR can detect unseen trap states by using failed demonstrations but primarily assigns higher rewards to states that frequently appear in successful trajectories and lower rewards to those that frequently appear in failed trajectories. This approach limits the agent’s ability to explore and find better solutions.
TW-CRL overcomes these limitations by enabling the agent to avoid trap states and efficiently discover paths to the goal. It effectively identifies both goal and trap states, encourages broader explorations, and allows the agent to uncover more efficient routes such as the shortcut path shown in Figure 4(c) and other alternative routes shown in Figure 4(d) which are not included in the expert demonstrations.
(a) TrapMaze-v1 map
(b) An expert demonstration
(c) A shortcut path
(d) An alternative path
Table 2: Average return and success rate under three goal reset settings. 1, 3, and any. #goal represent for the number of grids where the goal state might be in. The first sub-rows show the average return, and the second sub-rows show the success rate.
# goal state(s) 1 3 Any
TW-CRL 241.85 ± 8.80 254.79 ± 3.55 195.75 ± 17.94
0.94 ± 43.22 0.94 ± 0.01 0.75 ± 0.10
SASR 196.28 ± 32.89 150.84 ± 50.95 122.78 ± 20.39
0.82 ± 50.95 0.56 ± 0.19 0.47 ± 0.08
GAIL 159.06 ± 25.53 59.79 ± 22.87 55.25 ± 21.67
0.78 ± 50.95 0.36 ± 0.15 0.23 ± 0.14
AIRL 160.22 ± 11.25 74.26 ± 20.81 57.55 ± 30.27
0.72 ± 50.95 0.28 ± 0.11 0.23 ± 0.15
MaxEnt 129.70 ± 18.50 73.24 ± 20.10 49.77 ± 33.81
0.46 ± 50.95 0.24 ± 0.08 0.29 ± 0.16
5.4 Generalization Analysis
We then evaluate TW-CRL’s ability to generalize beyond the goal configurations provided in the original demonstrations in the TrapMaze-v1 environment, where the goal is randomly placed within one or several grid cells. Specifically, we provide demonstrations where the goal is initialized within a single grid cell, but during training and testing, the goal is allowed to be initialized within 1, 3, or any grid cells in the maze. Table 2 reports the average returns and success rates for each configuration. While all methods experience some drop in performance as the distribution of possible goal states broadens, TW-CRL remains consistently superior to the baselines, suggesting that its learned reward function can adapt successfully to unseen goal locations. Further results and details can be found in Appendix A and Appendix B.
5.5 Ablation Study
Effect of Time-Weighted function Figure 5(a) and Figure 5(b) (left) compare TW-CRL with time-weighted rewards (blue) to a simpler approach using constant rewards of $+1$ for success states and $-1$ for failure states (purple). Results show that time-weighting accelerates convergence and achieves higher final returns in both TrapMaze-v1 and TrapMaze-v2. The advantage is even clearer in TrapMaze-v2, where the time-weighted approach significantly outperforms the non-time-weighted version and vanilla TD3. These findings suggest that time weighting is particularly beneficial in complex tasks.
(a) TrapMaze-v1
(b) TrapMaze-v2
6 Conclusion
In this paper, we propose TW-CRL, an inverse reinforcement learning approach that infers reward functions by leveraging temporal information from both successful and failed demonstrations. By incorporating the Time-Weighted function and Contrastive Reward Learning, TW-CRL learns an accurate and informative reward function, enabling the agent to efficiently locate goals while avoiding repeated failures, thereby enhancing learning efficiency and increasing task success rates. Additionally, TW-CRL enhances exploration beyond expert demonstrations by discovering alternative high-value paths or shortcuts that were not explicitly demonstrated. Empirical evaluations 0demonstrate that TW-CRL outperforms baseline methods, showcasing its effectiveness in enhancing agent performance and learning efficiency in complex tasks.
Impact Statement
This paper presents work whose goal is to advance the field of Inverse Reinforcement Learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
Appendix A Experiment Environments
(a) U-Maze
(b) TrapMaze-v1
(c) TrapMaze-v2
(d) PandaReach
(e) PandaPush
@@ Appendix A Experiment Environments
We customized several environments based on the D4RL (Fu et al., 2020) PointMaze environment. In these custom environments, the agent’s actions correspond to displacement in the x and y directions, instead of applying forces in the x and y axes. We created a set of environments called TrapMaze. In these environments, several traps can cause the agent to get stuck. Once the agent is trapped, it cannot escape regardless of the actions it takes. In the classic U-Maze environment, the goal is represented by a red circle on the map, while the agent is shown as a green circle. The maximum number of timesteps allowed in this environment is 100. In the TrapMaze environments, blue circles represent the traps, the red circle represents the goal, and the green circle represents the agent. The positions of the traps can be configured using the character ‘t’ in the map code. We designed two types of TrapMaze environments for our experiments: TrapMaze-v1,which includes a gap in the solid wall that could potentially serve as a shortcut for the agent, and TrapMaze-v2. a more challenging version of TrapMaze-v1, where the shortcut is removed. In this version, if the agent’s reward is based solely on the distance to the goal, it will struggle to perform well. Instead, the agent needs to learn the actual maze layout to successfully reach the goal. Both TrapMaze environments have the maximum number of timesteps set to 300.
For the robotic manipulation benchmarks, we used the PandaReach and PandaPush environments from panda-gym (Gallouédec et al., 2021). All experimental environments are shown in Figure 6. In our experiments, we also used the PandaPush environment. Due to limited computational resources and to obtain results efficiently, we fixed the object’s position at (0.0, 0.0, 0.02). The goal was restricted to 12 fixed positions, with the goal location for each episode randomly selected from the following set: (0.1, 0., 0.02), (0.12, 0., 0.02), (0.15, 0., 0.02), (0.2, 0., 0.02), (0.1, -0.05, 0.02), (0.12, -0.05, 0.02), (0.15, -0.05, 0.02), (0.2, -0.05, 0.02), (0.1, 0.05, 0.02), (0.12, 0.05, 0.02), (0.15, 0.05, 0.02), (0.2, 0.05, 0.02).
@@
In the PointMaze environments, the goal is randomly initialized within specific grid cells configured by the maze map code. However, it cannot be placed in the same grid cell as the agent’s initial position or any traps. The grid cells where the goal can be placed are specified using the character ‘g’ in the maze map code. In Section 5.4, we evaluate the performance of TW-CRL under different configurations of initial goal grid placements. However, the provided demonstrations are only conducted in an environment where the goal is generated in a single grid cell, as shown in Figure 7(a). Figure 7 illustrates these configurations, where the goal can be randomly initialized within the areas shaded in white. Specifically, Figure 7(a) shows that the goal is initialized in a single specific grid cell, Figure 7(b) shows the goal initialized in three specific grid cells, and Figure 7(c) demonstrates the goal initialized in 13 specific grid cells, meaning it can appear in any grid cell within the maze.
Appendix B Additional Results
We provide additional results from our experiments here. Figure 8 shows the Training and success rate curves on TrapMaze-v1 under three different configurations mentioned in Section 5.4.
(a) Initialization in 1 grid
(b) Initialization in 3 grids
(c) Initialization in any grid
Figure 8: Training and success rate curves on TrapMaze-v1 under three different configurations. The solid curves represent the mean, and the shaded regions indicate the standard deviation over five runs. The first row shows the average return curves for each setting, while the second row shows the success rate curves for each environment.
Appendix C Experiment Details
This section describes the experimental setup and hyperparameters. In our experiments, TW-CRL is trained based on the TD3 reinforcement learning algorithm. The hyperparameters used are listed in Table 3.
@@ Appendix C Experiment Details
Time-Weighted hyperparameter ($k$) 1

### v2

@@ C.1 Experiment Environments
(d) MountainCar-v0
(e) HumanStand-v5
(f) Ant-v5
(g) PandaReach-v3
(h) PandaPush-v3
@@
We customized several environments based on the D4RL [8] PointMaze environment. In these custom environments, the agent’s actions correspond to displacement in the x and y directions, instead of applying forces in the x and y axes. We created a set of environments called TrapMaze. In these environments, several traps can cause the agent to get stuck. Once the agent is trapped, it cannot escape regardless of the actions it takes. In the classic U-Maze environment, the goal is represented by a red circle on the map, while the agent is shown as a green circle. The maximum number of timesteps allowed in this environment is 100. In the TrapMaze environments, blue circles represent the traps, the red circle represents the goal, and the green circle represents the agent. We designed two types of TrapMaze environments for our experiments: TrapMaze-v1,which includes a gap in the solid wall that could potentially serve as a shortcut for the agent, and TrapMaze-v2. a more challenging version of TrapMaze-v1, where the shortcut is removed. In this version, if the agent’s reward is based solely on the distance to the goal, it will struggle to perform well. Instead, the agent needs to learn the actual maze layout to successfully reach the goal. Both TrapMaze environments have the maximum number of timesteps set to 300.
For the robotic manipulation benchmarks, we use the PandaReach and PandaPush environments from panda-gym [11]. All experimental environments are shown in Figure 7. In our experiments, we also use the PandaPush environment. Due to limited computational resources and to obtain results efficiently, we fixed the object’s position at (0.0, 0.0, 0.02). The goal was restricted to 12 fixed positions, with the goal location for each episode randomly selected from the following set: (0.1, 0., 0.02), (0.12, 0., 0.02), (0.15, 0., 0.02), (0.2, 0., 0.02), (0.1, -0.05, 0.02), (0.12, -0.05, 0.02), (0.15, -0.05, 0.02), (0.2, -0.05, 0.02), (0.1, 0.05, 0.02), (0.12, 0.05, 0.02), (0.15, 0.05, 0.02), (0.2, 0.05, 0.02).
@@
In the PointMaze environments, the goal is randomly initialized within specific grid cells. It cannot be placed in the same grid cell as the agent’s initial position or any traps. In Section 5.3, we evaluate the performance of TW-CRL under different configurations of initial goal grid placements. In these experiments, the provided demonstrations are only conducted in an environment where the goal is generated in a single grid cell, as shown in Figure 8(a). Figure 8 illustrates these configurations, where the goal can be randomly initialized within the areas shaded in white. Specifically, Figure 8(a) shows that the goal is initialized in a single specific grid cell, Figure 8(b) shows the goal initialized in three specific grid cells, and Figure 8(c) demonstrates the goal initialized in 13 specific grid cells, meaning it can appear in any grid cell within the maze.
C.2 Experiment Details
This section describes the experimental setup and hyperparameters. In our experiments, TW-CRL is trained based on TD3. The hyperparameters used are listed in Table 4. One thing to note is that our experiment plots cover only a limited range of timesteps. As a result, the end points of the curves in the figures differ from the values in the results table, which report the fully converged results. We limit the plotted range to make differences in trend between methods more visible and to help readers compare them more easily.
@@ C.2 Experiment Details
MERIT
Hidden layers in discriminator network [256, 256, 256]
Discriminator learning rate $31e-$ 3
meta regularization hyperparameter ($\lambda_{m}eta$) $0.1$
ICRL
Hidden layers in discriminator network [256, 256, 256]
Discriminator learning rate $31e-$ 3
@@
Table 4: Detailed hyperparameters for each environment.
Hyperparameters Value
U-Maze
Time-Weighted Hyperparameter ($\alpha$) $2$
Horizon ($T$) $300$
TrapMaze-v1
Time-Weighted Hyperparameter ($\alpha$) $2$
Horizon ($T$) $300$
TrapMaze-v2
Time-Weighted Hyperparameter ($\alpha$) $2$
Horizon ($T$) $300$
MountainCar
Time-Weighted Hyperparameter ($\alpha$) $1$
Horizon ($T$) $999$
HumanStand
Time-Weighted Hyperparameter ($\alpha$) $0.1$
Horizon ($T$) $1000$
Ant
Time-Weighted Hyperparameter ($\alpha$) $0.1$
Horizon ($T$) $1000$
PandaReach
Time-Weighted Hyperparameter ($\alpha$) $2$
Horizon ($T$) $50$
PandaPush
Time-Weighted Hyperparameter ($\alpha$) $2$
Horizon ($T$) $100$
C.3 Hardware Information
We conducted our experiments on a GPU server. On 1 Nvidia L40 GPU and 8 Intel Xeon Platinum 8462Y CPUs, each run of TW-CRL takes around 4 to 24 hours depending on the complexity of the environment.
C.4 License Information
Our paper uses existing open-source libraries for our environments and baseline methods. All assets are under permissive licenses such as MIT and CC-BY-4.0, and we have made proper citations.
Appendix D Additional Results
D.1 Additional Results for Generalization
We provide additional results from our generalization experiments here. Figure 9 shows the average episode reward and success rate curves on TrapMaze-v1 under three different configurations mentioned in Section 5.3. Table 2 shows the average episode reward of other tested baseline methods not shown in Table 2 due to page limit.
Table 5: Additional average return under three goal reset settings. 1, 3, and any. #goal represents the number of grids where the goal state might be in.
# goal state(s) TW-CRL TD3 MaxEnt AIRL GAIL
1 253.12 ± 9.46 196.28 ± 32.89 129.70 ± 18.50 160.22± 11.25 159.06 ± 25.53
3 240.48 ± 4.44 168.58 ± 27.86 73.24 ± 20.10 74.26 ± 20.81 59.79 ± 22.87
Any 221.33 ± 11.58 168.38 ± 40.06 49.77 ± 33.81 57.55 ± 30.27 55.25 ± 21.67
(a) Initialization in 1 grid
(b) Initialization in 3 grids
(c) Initialization in any grid
Figure 9: Training and success rate curves on TrapMaze-v1 under three different configurations. The solid curves represent the mean, and the shaded regions indicate the standard deviation over five runs. The first row shows the average return curves for each setting, while the second row shows the
D.2 Additional Results for Reward Function Visualizations
We provide additional results of visualization of reward function in TrapMaze-v1 environment in Figure 10.
Figure 10: Visualization of the reward function in TrapMaze-v1.
Appendix E Limitation and Future Work
In this paper, we introduce TW-CRL as a new IRL method that uses both successful and failed demonstrations along with the temporal information of each state to learn a better reward function and train an agent’s policy. In its current form, TW-CRL works best on tasks where a series of actions unfolds over time, such as navigation problems. It is not easily applicable by design for tasks that require highly repetitive or back-and-forth actions, like Humanoid or the HalfCheetah. Its performance also depends a lot on how the Time-Weighted function and the success threshold are set. With careful tuning of these hyperparameters, TW-CRL can be adapted to other environments, but this tuning must be done separately for each new task.
In future work, we will explore different forms of the time-weight function and test how they affect learning. We will also investigate using TW-CRL for reward shaping by adding the learned reward to the environment’s own reward, which may speed up training on tasks that already provide clear feedback. Finally, we aim to develop simple methods for automatically selecting the Time-Weighted function and success threshold so that TW-CRL can be applied more robustly across a wider range of environments.